T1DISASM(1) T1DISASM(1) NAME t1disasm−disassembletype-1font SYNOPSIS t1disasm[input[output]] DESCRIPTION t1disasmdisassemblesAdobeType1fontprogramsineitherPFA(hexadecimal)orPFB (binary)formatsintohuman-readableform.Ifthefileoutputisnotspecifiedoutputgoestothe standardoutput.Ifthefileinputisnotspecifiedinputcomesfromthestandardinput. t1disasmperformseexecandcharstringdecryptionasspecifiedinthe‘‘blackbook’’,AdobeType 1FontFormat.Additionally,thecharstringbinarytokensareexpandedintohuman-readabletext form,usingthenamesgivenintheblackbookandlaterdocumentsdescribingType2opcodes. EXAMPLES %t1disasmUtopia-Regular.pfbUtopia-Regular.raw %t1disasmUtopia-Regular.pfaUtopia-Regular.raw InSubrsentriesinUtopia-Regular.rawwilllooklike dup5{ 8111vstem -12128hstem 707-20hstem return }| andtheCharStringsentrieslike /exclam{ 58242hsbw 6callsubr 54callsubr 63707rmoveto -540-5-224-45rrcurveto 40-431rlineto 29hlineto 42431rlineto 445-522-550rrcurveto closepath 64callsubr -719vmoveto 243callsubr endchar }|- SEEALSO t1asm(1),t1ascii(1),t1binary(1),t1unmac(1),t1mac(1) AdobeType1FontFormatisavailablefreefromAdobeasaPDFfile:http://partners.adobe.com/asn/developer/PDFS/TN/T1SPEC.PDF TheType2CharstringFormat,alsoavailablefromAdobeasaPDFfile,describesthenewerType 2operators,whicharealsousedinsomemultiple-masterType1fontslikeAdobeJensonand Kepler:http://partners.adobe.com/asn/developer/PDFS/TN/5177.Type2.pdf AUTHORS LeeHetherington(ilh@lcs.mit.edu) EddieKohler(eddietwo@lcs.mit.edu) 5Mar1998 1